OpenStudy (anonymous):

The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

A prism has identical top and bottom surfaces and it has lateral surface (side surface). Total surface area = 2 * area of base + perimeter * height

OpenStudy (anonymous):

I got 980...?

pooja195 (pooja195):

Bottom area = 55 Top area will be the same as the bottom. Therefore, 55 Lateral or side surface area = perimeter of base * height = 30 * 7 = 210 Total surface area = 210 + 55 + 55

pooja195 (pooja195):

that should get you the answer

OpenStudy (anonymous):

B

pooja195 (pooja195):

yes its B good job :D \(\huge\color{blue}{WELCOME~TO~OPENSTUDY!!!}\)

OpenStudy (anonymous):

THANK YOU!!

pooja195 (pooja195):

Your welcome :D